A. Balance sheet effects: the inventory costs are closer to current costs under FIFO than under LIFO.
b. Income statement effects: in addition to the effects on net income in (13) above, LIFO enables the company to avoid reporting paper or phantom profit as economic gain.
c. Tax effects: in a period of inflation LIFO results in the lowest income taxes.

Which method do you prefer in term of showing higher net income? why
Answer #1

For showing higher net income in a business environment where the inventory cost keeps increasing (inflation), I would choose to value inventory using FIFO method.

The FIFO method evaluates inventory on the basis of the concept "whatever comes in first, goes out first". So the items remaining inventory will be closer to the market price unlike in the LIFO method in which "whatever comes in first, goes out last". In a business environment where there is inflation, the cost of purchases increases with every purchase. Under FIFO, the items in inventory purchased first (with lower cost), will be sold first. So the cost of goods sold will be less as compared to LIFO method because under LIFO method, the items purchased last (with higher cost) are sold first thus resulting in a higher cost of goods sold. Let us understand the difference between these two inventory valuation methods with the help of an example in a perpetual inventory system:

Inventory Records
Particulars $
Opening inventory(1/1/2018), 20 units at $10 per unit 200
Purchases (1/10/2018), 50 units at $12 per unit 600
Sales (1/20/2018), 60 units at $20 per unit 1200
Purchases (1/29/2018), 20 units at $13 per unit 260

as we can see the purchase price of the units increases everytime a new purchase is made, depicting a situation of inflation.

Cost of goods sold under:

  1. FIFO method= (20*10) + (40*12)= $680
  2. LIFO method= (50*12) + (10*10)= $700

As can be seen clearly, under FIFO method, the cost of goods sold is lower and as a result, the net income will be higher as compared to LIFO method. Therefore, to show higher net income FIFO method should be chosen in an inflating environment.

However, had inventory cost decreased with every purchase (negative inflation), LIFO method would have given lower cost of goods sold and a higher net income.